OpenStudy (anonymous):

Y=2x-1 find the slope of a line perpendicular to each given line

OpenStudy (anonymous):

Plus you are posting the questions so fast that you can't possibly be reading and understanding any of the answers that you are getting.

OpenStudy (anonymous):

omg!im not gonna get these brief explanations anyway

OpenStudy (anonymous):

Well if you would follow up with some kind of response rather than just posting more problems you might get a longer explanation. What is confusing to you about the problems you posted?

OpenStudy (anonymous):

everything.

OpenStudy (anonymous):

From the equation you posted, can you identify the slope of the line?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If you have x = 1, what is y?

OpenStudy (anonymous):

the general form of a line is y=mx + b where m is the slope, and b is a constant

OpenStudy (anonymous):

This might help: http://tutorial.math.lamar.edu/Classes/Alg/Lines.aspx

OpenStudy (anonymous):

ok i knw the y=mx+b thing.but is it 0?

OpenStudy (anonymous):

If x = 1 then y is 0 yes. Now if x increases by 1 more (goes to 2) then how much does y increase by?

OpenStudy (anonymous):

no, the line is Y=2x-1 y=mx+b so the m is 2, and the b is -1

OpenStudy (anonymous):

oooo!ok thats simple

OpenStudy (anonymous):

Err wait, sorry. No if x = 1 y = 1.

OpenStudy (anonymous):

The slope of the line is "How much does y increase (or decrease) as x increases

OpenStudy (anonymous):

wait now imm confused again.m=2?

OpenStudy (anonymous):

So if x goes up by 1, y will go up by 2 more.

OpenStudy (anonymous):

actually you have another way to find the slope. first you need to find two point that's passed by the line. just plug x=0 -> y=2(0)-1 -> y=-1 then plug y=o -> 0=2x-1 -> x=1/2 now you have two points (0,-1), this is the (x1,y1). and (1/2 ,0), the (x2,y2) the formula to find slope is : m= (y2-y1)/(x2-x1) then just plug the point, you'll get m

OpenStudy (anonymous):

If x = 1, y = 2(1) -1 = 2-1 = 1. If x = 2, y = 2(2) - 1 = 4-1 = 3. If x = 3, y = 2(3) -1 = 6 -1 = 5. So we can see that when x increases by 1, y increases by 2. So the slope of the line is 2.

OpenStudy (anonymous):

Once you have determined the slope of the line, you can find a perpendicular line by inverting it (divide 1 by it) and make it the opposite sign.

OpenStudy (anonymous):

Did you finish this one? or the other 6 you posted?